OpenStudy (anonymous):

what is the d'(y)= (y^2)^2+(y-3)^2?? any help?

OpenStudy (anonymous):

Are you asking for the first derivative with respect to y? If so, then you can do this many ways, the easiest is to simplify first. y^2^2=y^4, (y-3)^2 = y^2-6y+9. Now just take the derivative of y^4+y^2-6y+9 with respect to y. This becomes 4y^3+2y-6

OpenStudy (anonymous):

how would i find what y is worth from that equation .. because thats what i had before and i dont know where im going wrong?>

OpenStudy (anonymous):

What do you mean find out what y is worth? Y is a variable, it does not have a set value.

OpenStudy (anonymous):

critical points i meant of the derivative

OpenStudy (anonymous):

If you want critical points of the original function, then set the derivative equal to zero.

OpenStudy (anonymous):

ya i know this ... but i cant find the right answer to it .. how would you do that when you have a third root of somthing ...

OpenStudy (anonymous):

Take factors of the first term and then take factors of the second term. After that, you have to do some sythetic division.

OpenStudy (anonymous):

sorry not second term i meant last term